OpenStudy (anonymous):

Steve has the following scores on the exams for the semester: 89,92,87. What will he need to make on the fourth exam to have a final average of a 92 or better? Write the answer as an inequality

OpenStudy (anonymous):

So, first find the average of 3 semester scores.. Tell me what is the sum of 89, 92 and 87???

OpenStudy (anonymous):

The average is 89.333

OpenStudy (anonymous):

Of what???

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Of the three semesters…

OpenStudy (anonymous):

See, now let the marks scored in 4th semester be x: do you get this: \[Average = \frac{89 + 92 + 87 + x}{4}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now you know that average = 92 and: \[89 + 92 + 87 = 89.3 \times 3\] This one you got or not??